OpenStudy (anonymous):

WILL GIVE A MEDAL can anybody tell me the equation for the ellipse of Earth's orbit?

OpenStudy (anonymous):

@aum

OpenStudy (aum):

\[ \frac{x^2}{a^2} + \frac{y^2}{b^2} = 1 \]For Earth, a = 1 AU (Astronomical Unit); b = 0.99993 AU; c = 0.0085 AU a^2 = 1; b^2 = 0.99986 \[ \frac{x^2}{1^2} + \frac{y^2}{0.99993^2} = 1 \]where the Sun is at one of the focal points. This is very nearly a circle. If you want the Sun to be the center of the ellipse instead of the focal point, then shift the above equation to the left (or right) by 'c' which is 0.0085 AU. http://jwilson.coe.uga.edu/EMAT6680Fa08/Broderick/essay1/essay1.html